# Hardness of Regular Expression Matching with Extensions

### v1

@@ 6 Conclusion
In this paper, we have studied the hardness of regular expression matching with three well-known extensions, backreference, intersection and complement. We have shown that (i) neither 1-rewb matching nor semi-ERE matching can be solved in $O(n^{2-\varepsilon}\operatorname{poly}(m))$ time for any $\varepsilon>0$ under the Orthogonal Vectors Conjecture (Theorems 1 and 2), and that (ii) ERE matching cannot be solved in $O(n^{\omega-\varepsilon}\operatorname{poly}(m))$ time for any $\varepsilon>0$ under the $k$-Clique Hypothesis and it cannot be solved combinatorially in $O(n^{3-\varepsilon}\operatorname{poly}(m))$ time for any $\varepsilon>0$ under the Combinatorial $k$-Clique Hypothesis (Theorem 3). As mentioned in the introduction, our results entail a number of interesting theoretical and practical consequences. For instance, our results provide evidence that unlike the case of lookaround, $O(nm)$-time matching algorithms for regular expressions extended with these features (in their full generality) are unlikely to exist, and that the ERE matching algorithm by Hopcroft and Ullman [23] was already optimal in the sense described in Section 1.2.
A possible direction for future work is to close the gap between the upper and lower bounds of 1-rewb matching and semi-ERE matching. Regarding upper bounds for 1-rewb matching, as mentioned in Section 2, Nogami and Terauchi [33] recently proposed an $O(n^{2}m^{2})$-time algorithm for a certain subclass of 1-rewbs. If the general 1-rewb matching can also be solved in the same time complexity (or up to a polynomial factor in $m$), Theorem 1 provides evidence under OVC that there can be no further significant improvement in $n$ unless we allow a superpolynomial blowup in $m$. However, generalizing their algorithm to all 1-rewbs does not seem straightforward, as the algorithm makes non-trivial uses of the special structure of the 1-rewb subclass that it considers.

### v2

@@ 6 Conclusion
In this paper, we have studied the hardness of regular expression matching with four well-known extensions, backreference, squaring, intersection and complement. We have shown that (i) neither 1-rewb matching, RSQ matching nor semi-ERE matching can be solved in $O(n^{2-\varepsilon}\operatorname{poly}(m))$ time for any $\varepsilon>0$ under the Orthogonal Vectors Conjecture (Theorems 1, 2 and 3), and that (ii) ERE matching cannot be solved in $O(n^{\omega-\varepsilon}\operatorname{poly}(m))$ time for any $\varepsilon>0$ under the $k$-Clique Hypothesis and it cannot be solved combinatorially in $O(n^{3-\varepsilon}\operatorname{poly}(m))$ time for any $\varepsilon>0$ under the Combinatorial $k$-Clique Hypothesis (Theorem 4). As mentioned in the introduction, our results entail a number of interesting theoretical and practical consequences. For instance, our results provide evidence that unlike the case of lookaround, $O(nm)$-time matching algorithms for regular expressions extended with these features (in their full generality) are unlikely to exist, and that the ERE matching algorithm by Hopcroft and Ullman [29] was already optimal in the sense described in Section 1.2.
A possible direction for future work is to close the gap between the upper and lower bounds of 1-rewb matching, RSQ matching and semi-ERE matching. Regarding upper bounds for 1-rewb matching, as mentioned in Section 2, Nogami and Terauchi [45] recently proposed an $O(n^{2}m^{2})$-time algorithm for a certain subclass of 1-rewbs. If the general 1-rewb matching can also be solved in the same time complexity (or up to a polynomial factor in $m$), Theorem 1 provides evidence under OVC that there can be no further significant improvement in $n$ unless we allow a superpolynomial blowup in $m$. However, generalizing their algorithm to all 1-rewbs does not seem straightforward, as the algorithm makes non-trivial uses of the special structure of the 1-rewb subclass that it considers. Another possible direction for future work is to prove lower bounds on the matching problem with bounded repetition of constant nesting depth. This would provide further insights on the hardness of bounded repetition in practice, where it is typically not deeply nested.